OpenStudy (anonymous):

Consider the function f(x) = 1/3x+ 2 Use function composition to show that f(x) and g(x) are inverses of each other.

hartnn (hartnn):

where is g(x) ? or u need to find it ?

OpenStudy (anonymous):

need to find it. @hartnn

hartnn (hartnn):

so let y= 1/3x+2 exchange x and y x=1/3y+2 can u find y in terms of x now ?

OpenStudy (anonymous):

is it 3x-6?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes. that is g(x). now find f(g(x))

OpenStudy (anonymous):

1/3 (3x-6)+2 = x-2+2=x ? @hartnn

OpenStudy (anonymous):

or is it 1/3(3x-6)+2=3(1/3x+2)-6 ?

hartnn (hartnn):

good work :) since f(g(x))=x u have verified that f and g are inverse of each other. also prove that g(f(x)) = x

OpenStudy (anonymous):

wait which one was correct? the first one i said or the second. ?

hartnn (hartnn):

first.

OpenStudy (anonymous):

ok. so now how do i graph this?

OpenStudy (anonymous):

I have to draw the graphs of f(x) and g(x) on the same coordinate plane.

hartnn (hartnn):

f(x)=y=1/3x+2 put x=0, y=? put y=0,x=?

OpenStudy (anonymous):

put x = 0, y= 1/3?

OpenStudy (anonymous):

Im not sure... @hartnn

hartnn (hartnn):

put x=0 y=1/3 x + 2 y=0+2 y=2. ok?

OpenStudy (anonymous):

ok so 0,2? and 0,-6?

OpenStudy (anonymous):

those are the coordinates?

hartnn (hartnn):

0,2 is correct (-6,0)

OpenStudy (anonymous):

ok so (0,2) and (-6,0)

OpenStudy (anonymous):

i draw this on the x-axis? or y axis?

hartnn (hartnn):

yup, plot those points. and join them by straight line..

hartnn (hartnn):

(0,2) will lie on y axis. (-6,0) will lie on x-axis

OpenStudy (anonymous):

the -6 gets a closed circle and the 2 is open right?

hartnn (hartnn):

circle ?

OpenStudy (anonymous):

never mind. that's for a different one :D THANKS FOR YOUR HELP!

OpenStudy (anonymous):

how do i eXPLAINE how my grapgh shows that the functions are inverses of each other?

hartnn (hartnn):

sorry for late reply. u can verify that f and g are inverses of each other by seeing that when f is reflected over line y=x , we get g. here's the plots: http://www.wolframalpha.com/input/?i=plot+x%2F3%2B2++and+3x-6+and+x

hartnn (hartnn):

u asked same question @lavalava did u get all 3 parts ?

hartnn (hartnn):

@Riza still any doubts ?

OpenStudy (anonymous):

@hartnn no sir thanks